OpenStudy (anonymous):

Graph the following linear system to determine the number of solutions. y = -3x + 6x = 2y = 12

OpenStudy (bahrom7893):

y = -3x + 6x 2y = 12 was this the original question?

OpenStudy (anonymous):

yess

OpenStudy (bahrom7893):

y = -3x+6x = 3x 2y = 12; y = 6 --------------- y=3x y=6, so: 6=3x; x=2. There's one solution: x=2, y=6

OpenStudy (anonymous):

Okaay. THANK U

OpenStudy (anonymous):

Graph the following linear system to determine the number of solutions. y = -x + 2 x + 5 = y